OpenStudy (anonymous):

Solve the systems of equations for p: p + 2/5q = 6 3/2p - q = 1 When I did this, I got p=8 and q=-5. I just want to check if I'm right. Please help XD

OpenStudy (loser66):

\[p +\frac{2}{5}q=6\\\frac{3}{2}p-q=1\]or\[p+\frac{2}{5q}=6\\\frac{3}{2p}-q=1\]which one?

OpenStudy (anonymous):

The first one. My bad.

OpenStudy (loser66):

nope, i plug your solution into the set and check, the first one is ok ,but the second one \[\frac{3}{2}8-(-5)=12+5=17~is~not=1\]so, yours is wrong

OpenStudy (anonymous):

So q is right, but p is wrong?

OpenStudy (loser66):

they HAVE to go TOGETHER, you cannot break them out.

OpenStudy (loser66):

redo until you get pair of them which satisfy both equations. Keep in mind, "PAIR" of them,

OpenStudy (anonymous):

Okay, thanks

OpenStudy (loser66):

yw